Suppose that you are given a $100 budget at work that can be spent only on two items: staplers and pens. If staplers cost $10 each and pens cost $2.50 each, then the opportunity cost of purchasing one stapler is: 10 pens 5 pens 04 pens. zero pens
Answer #1

BUDGET $100

STAPLERS COST $ 10

PENS COST $2.50

OPPORTUNITY COST OF PURCHASING ONE STAPLER

TOTAL NUMBER OF STAPLERS = 100/10

TOTAL NUMBER OF STAPLERS = 10

TOTAL NUMBER OF PENS = 100/2.50

TOTAL NUMBER OF PENS = 40

OPPORTUNITY COST OF PURCHASING ONE STAPLER = TOTAL QUANTITY OF PENS/TOTAL QUANTITY OF STAPLERS

OPPORTUNITY COST = 40/10

OPPORTUNITY COST = 4

CORRECT ANSWER IS 4 PENS